please help!!! simplify showing your steps 1 over 4 (1.4) + (-4+4)
: please help!!! simplify showing your steps 1 over 4 (1.4) + (-4+4)

Answer by Alan3354(1941)

please help!!! simplify showing your steps 1 over 4 (1.4) + (-4+4)
----------------------
Is it 1/(4*1.4)+(-4+4)) ??
That's 1/5.6 + 0
= 1/5.6
= 10/56
= 5/28
= 0.17857142857...